OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (anonymous):

the substitution method involves substituting one variable for another

OpenStudy (sdfgsdfgs):

if u add 4y to the left and right of the eqn x - 4y =1, what will u get?

OpenStudy (anonymous):

My bro sdfgsdfgs just wanted u to perfectly understand the question before applying this method. He's old school like me!

OpenStudy (anonymous):

Sorry my brother, @sdfgsdfgs now takes over

OpenStudy (sdfgsdfgs):

@BPDlkeme234 no need to apologize - u actually posted be4 i did ;) @kaylamarie081 would u like to work thro the prob or u waiting 4 someone to give u the ans?

OpenStudy (anonymous):

i just need help through it all. i dont understand what to do.

OpenStudy (sdfgsdfgs):

@kaylamarie081 OK good :) lets work thro it together! if u add 4y to the left and right of the eqn x - 4y =1, what will u get?

OpenStudy (anonymous):

5y

OpenStudy (anonymous):

Kaylamarie, thats a nice name!

OpenStudy (sdfgsdfgs):

Try again plz. it should be like this: x - 4y + 4y = 1 + 4y

OpenStudy (anonymous):

Just doing some rough work here. If you were to use the substitution method you might say: x= 1+4y

OpenStudy (sdfgsdfgs):

So x = ?

OpenStudy (anonymous):

you micht substitute the value of this in equation 1 to get: 2(1+4y) -9y =1

OpenStudy (anonymous):

which might give you the following: 2+ 8y -9y = 1

OpenStudy (anonymous):

From which you might find: 2-y =1 or to put it another way: -y = 1 -2 i.e y = -1+2

OpenStudy (anonymous):

From there it is a snows drop away to find the answer, that y = 1.

OpenStudy (anonymous):

Now all you need to do is substitute this value into either equation (it really doesnt't matter) in order to find the corresponding value for x

OpenStudy (anonymous):

if any medals should go for this solution is should be to @sdfgsdfgs who showed me the solution

OpenStudy (sdfgsdfgs):

@BPDlkeme234 bro u are way too modest! i wuld give u another medal if i could!

OpenStudy (anonymous):

there he is again, @sdfgsdfgs cant accepte he provided the solution!

OpenStudy (anonymous):

He looked at the bakance of the equations and he knew there was something wrong

OpenStudy (anonymous):

And he stepped right in there to correct it, while at the same time giving me his personal thoughts about this problem

OpenStudy (anonymous):

All I did was reproduce the stuff he said

OpenStudy (anonymous):

I picked up a medal for that, thanks @sdfgsdfg